# LLM Inference Unveiled: Survey and Roofline Model Insights

### 11 Mar 2024 v4

@@ 6 Hardware Optimization
Designing hardware to efficiently support inference for LLMs is a challenging task due to the significant differences in arithmetic intensity compared to previous tasks.  [^6] Arithmetic intensity refers to the ratio of arithmetic operations to memory access, which has been described in Roofline model (Section 2.3). This challenge arises primarily from one key factors: Low Arithmetic Intensity during Decoding: The model performs relatively few arithmetic operations compared to the large amount of memory access required (weight and KV cache), leading to a low arithmetic intensity. Variability in Arithmetic Intensity: The arithmetic intensity can vary greatly across prefill/decode stages, batch sizes, or sequence lengths. This variability complicates the hardware design process, as different stages or configurations may require different optimization strategies.
To address these challenges, careful consideration and optimization of hardware designs are necessary. In this section, we will survey and analyze various hardware optimizations specifically tailored for efficient LLM inference, with a focus on addressing the issues related to arithmetic intensity.
@@ 6.1 Spatial Architecture
In one case, Groq employs their LPU [Abts et al., , 2022] to create a spatial system for LLM inference. This system achieves a remarkable speed of over 300 tokens per second per user on the Llama-2-13b model [Groq, , 2023]. Another example is Graphcore’s Intelligence Processing Unit (IPU), which is another type of spatial architecture that efficiently executes LLMs [Graphcore, , 2024].
6.2 In-Memory and Near-Memory Computing
@@
The decoding phase of LLM inference experiences the so-called ”Memory Wall” problem, primarily due to its low arithmetic intensity. This issue is not new; the computer architecture community has been struggling with the ”Memory Wall” problem for several decades. Among various potential solutions, in-memory or near-memory processing techniques have garnered significant interest in recent years. By placing compute units into DRAM memory chips, we can leverage the much higher internal memory bandwidth and reduce the data movement overhead. This technique is called Near-Memory Computing (NMC). Samsung’s HBM PIM and Sk-hynix’s GDDR PIM are examples of products that use this technique. However, NMC currently has limited computational capacity, which makes it less suitable for the prefill stage and large batch sizes.
Process-in-Memory (PIM) is another emerging technology that has potential to address the challenges associated with the inference of LLMs and other data-intensive workloads. Traditional von Neumann architectures, where data is transferred between the processor and memory, can become a bottleneck for these computationally intensive tasks. PIM architectures integrate processing units directly into memory chips or memory modules. This allows computation to be performed within the memory itself, reducing the need for data movement between the processor and memory. With LLMs, we can store weights in the memory and input activation data to execute multiplication-and-accumulation (MAC) in the linear layer. This reduces latency and energy consumption associated with data transfer, resulting in improved performance.
Although PIM and NMC are still in early stages and face challenges in programming models, memory organization, and architectural design, they show promise in solve the problem of ”Memory Wall”, enhancing the performance and energy efficiency of LLM inference.

### 15 Mar 2024 v5

@@ 6 Hardware Optimization
Designing hardware to efficiently support inference for LLMs is a challenging task due to the varying arithmetic intensity [^6] Arithmetic intensity refers to the ratio of arithmetic operations to memory access, which has been described in Roofline model (Section 2.3). under different inference stages and workload conditions. Specifically, the prefill stage usually leverages GEMM operators to process the batched tokens, which exhibits high arithmetic intensity. On the contrary, the decoding stage calculates output tokens one at a time, which necessitates the use of either GEMV operators or lean GEMM operators to process the attention and FFN layers. These operators are characterized by low arithmetic intensity.
Furthermore, the arithmetic intensity can exhibit substantial variation depending on the batch sizes and sequence lengths. For instance, a large batch size could significantly alter the arithmetic intensity, and a long sequence length may increase the memory access overhead of KV-cache reading in each decoding step. This variability introduces additional complexity into the hardware design process, as different stages or configurations may necessitate distinct optimization strategies. Hence, it’s crucial to consider these factors when designing hardware to ensure efficient performance across a wide range of scenarios.
Considering these challenges, careful consideration and optimization of hardware designs are necessary. In this section, we will survey and analyze various hardware optimizations tailored for efficient LLM inference, with a focus on addressing the issues related to varying arithmetic intensity.
@@ 6.1 Spatial Architecture
In one case, Groq employs their LPU [Abts et al., , 2022] to create a spatial system for LLM inference. This system achieves a remarkable speed of over 300 tokens per second per user on the Llama-2-70b model [Groq, , 2023]. Another example is Graphcore’s Intelligence Processing Unit (IPU), which is another type of spatial architecture that efficiently executes LLMs [Graphcore, , 2024].
6.2 Processing in Memory
@@
The decoding phase of LLM inference experiences the so-called ”Memory Wall” problem, primarily due to its low arithmetic intensity. This issue is not new, the computer architecture community has been struggling with the ”Memory Wall” problem for several decades. Among various potential solutions, processing in-memory techniques (PIM) have garnered significant interest in recent years. By placing compute units directly in memory chips, we can leverage the much higher internal memory bandwidth and reduce the data movement overhead between memory and CPU/GPU cores.
In recent years, DRAM-based PIM has entered the commercialization phase, which can potentially mitigate the memory bandwidth bottleneck of LLM inference. As listed in Table 2, UPMEM’s PIM-DIMM Devaux, 2019a is the first commercialized DRAM-PIM product, which places general-purpose RISC cores in DDR4-DIMMs. However, this product was not intended for deep-learning applications, thus the peak bandwidth and throughput can hardly meet the requirements of LLM inference. Compared to UPMEM’s PIM-DIMM, Samsung proposes to place MAC units into HBM memory, achieving 2TB/s of internal memory bandwidth, which is much higher than that of traditional HBM2 (307GB/s per cube) memory. Since the processing units are tailored for deep learning workloads, the peak compute throughput of HBM-PIM can reach 1.2TFLOPS. In other words, HBM-PIM is suitable for accelerating operators with 1-2 Ops/Byte of arithmetic intensity.
Choi et al has proposed to accelerate the KV-Cache processing using HBM-PIM, which has low arithmetic intensity in batched LLM inference. According to their evaluation Choi et al., [2023], A GPU+HBM-PIM system for LLM inference can achieve 3.24$\times$ speedup over traditional monolithic GPU system.
Similar to Samsung’s HBM-PIM, SK-hynix has also proposed a GDDR6-based PIM accelerator Kwon et al., [2022] called AiM. As shown in Table 2, the compute units of AiM adopts the BF16 data format, which is more efficient for deep-learning acceleration. With optimized MAC units, AiM offers 1TFLOPS per chip of compute capacity, while the peak bandwidth is 1TB/s per chip. Although AiM has not reported its performance on LLM, it can achieve up to 10$\times$ speed improvement compared to GPU+HBM2 system on LSTM tasks.
Note that, although DRAM-based PIM techniques have demonstrated promising potential of accelerating memory-intensive operators in LLM inference, there are still some limitations that should be addressed in the future.
  • Limited Computation Power. A key restriction of DRAM-PIMs in accelerating LLMs stems from their constrained computational capabilities. DRAM-PIMs utilize compute units crafted using the DRAM process, resulting in transistors that are 3 times slower and logic density several times lower compared to CMOS in the same technology node Devaux, 2019b . Even worse, DRAM chips usually have fewer metal layers, leading to a lower routing density at the same time. Due to these technical constraints, DRAM-PIMs can hardly incorporate powerful compute units. As a consequence, DRAM-PIMs is only suitable for small-batch inference or KV-cache processing. For computation-intensive large-batch inference, a powerful host is still necessary.
  • Capacity Constraints. Another significant limitation of DRAM-PIMs is their restricted capacity. Given that DRAM-PIMs allocate some memory capacity to construct the computation units, the total memory capacity is typically 50% less than that of standard memory Kwon et al., [2021]. For LLM applications that necessitate substantial memory capacity to accommodate the weights and KV-cache, DRAM-PIMs may encounter capacity-related challenges.
  • Inadequate Inter-PIM Communication. In addition to the constraints imposed by computation power and capacity, another limitation of DRAM-PIMs is their suboptimal inter-PIM communication capability. Given that there are distributed computing units located near each DRAM bank, data aggregation and computation synchronization among these units are unavoidable. However, DRAM-PIMs lack robust interconnects Zhou et al., 2023c , Jonatan et al., [2024], and they typically depend on the host CPU/GPU for data exchange between PIM units. This reliance can lead to inefficiencies in the system. Therefore, to enhance LLM inference, future iterations of DRAM-PIMs should aim to improve their inter-PIM communication abilities.
Table 2: Comparison of Commodity DRAM NMC Products
Product PIM-DIMM HBM-PIM AiM
Technique DDR4 HBM2 GDDR6
PIM Units RISC Cores FP16 MAC BF16 MAC
Peak Bandwidth 80.4 GB/s per DIMM 2 TB/s per cube 1 TB/s per chip
Peak Throughput 43.8 GOP/s per DIMM 1.2 TFLOPS 1 TFLOPS